# Crossing Boundaries: Leveraging Semantic Divergences to Explore Cultural Novelty in Cooking Recipes

### v1

@@ Abstract
Keywords First keyword  $\cdot$ Second keyword  $\cdot$ More
@@ 1 Introduction
In the context of NLP and AI, the emergence of generative large language models (LLMs) has spurred research into cultural diversity and awareness, primarily through external tasks involving prompting (Bhatt and Diaz,, 2024; Li et al.,, 2024). Among such approaches, researchers try to measure LLMs personality traits, values and cultural perspective through self-assessment tests and controllable prompts (Kovač et al.,, 2023). Recent work using the WVS Inglehart–Welzel cultural map (,, WVS) proposed interpretable cultural distance metrics to evaluate LLM outputs (Tao et al.,, 2024). While this approach highlights cultural representation differences, its reliance on specific questionnaires limits generalizability across tasks and datasets. In this paper, we propose five new cultural novelty detection metrics based on probability distribution divergences, an area of interest in out-of-distribution detection (Yang et al.,, 2024), which exhibits promising properties, such as entropic uncertainty, when applied to LLMs (Ahmadian et al.,, 2024). In sociology and management, divergence measures have been widely used to analyze cultural evolution, legal systems, societal norms (Klingenstein et al.,, 2014), and community values (Gallagher et al.,, 2018). This work bridges the gap between cultural novelty and distance in sociology and the challenges of cultural representation and diversity in AI. Leveraging divergence-based novelty detection, we propose new metrics to measure cultural differences through text analysis, focusing on cultural proxies like food. Our contributions are as follows:
@@ 4.1 Dataset
To build our dataset on cultural novelty, we used RecipeNLG (Bień et al.,, 2020). The dataset notably comprises the title, a recipe description, and a list of ingredients of more than 2M recipes. To extract our culturally aware knowledge bases and variations, we kept the recipes where we could identify a country in the title. We ended up with 135 385 recipes from 173 different countries. Then, we defined product names as typical dishes, for example, "Lasagna," "Couscous," or "Beef Curry," which were associated with a country as our cultural variable. Using the recipe title to match the proper dishes, we extracted each cultural product’s associated recipe description to establish our knowledge space. We also identified variations of these dishes as recipe descriptions from different countries. Figure 2 presents an illustration of an image of recipes extracted with this approach for the couscous recipes, where the knowledge space is composed of 167 Moroccan recipes and 426 cultural variations coming from 35 different countries of all regions of the world.
@@
We identified 500 dishes, associated with 54 various countries from different geographical regions, forming our cultural product pairs (product, country). We used the title to match a recipe and extracted the description to create our knowledge base. Each dish’s knowledge space comprises an average of 60.98 recipe descriptions and is paired with 208.96 recipe descriptions of new variants coming from 19.28 different countries, for a total of 104,710 recipes variations. The specific protocol for data collection, ensuring respect for several conditions presented and detailed statistics on the dataset are provided in the appendix A.
@@ 4.2 Experiment and Evaluation
  • We use the Inglehart–Welzel Cultural Map (,, WVS) provided by the World Value Survey. This global cultural map scores society on two dimensions: (1) Traditional values versus Secular-rational values and (2) Survival values versus Self-expression values. Following the work in (Tao et al.,, 2024), this two-dimensional measurement of culture allows us to estimate the Euclidean distance between countries studied.
  • We use the linguistic distance dataset [^3] http://dow.net.au/?page_id=32. We used the data collected from 2015 which followed the research protocol from (Gordon Jr,, 2005). This distance represents the distance between the two closest major languages for each pair of countries.
@@
The dataset and the code are available on our Github page [^5] https://github.com/fcarichon/Cultural_Novelty/tree/main.
@@ 5 Results
Figure 4: Correlations between our 5 novelty metrics and 4 distinct cultural distances. The bars’ height highlights the strength of the coefficient, while the color intensity is the strength of its p-value; opaque colors indicate statistically significant correlations, while transparent colors indicate non-significative ones.
@@ 6 Discussion
This present study took an interest in the link between textual novelty detection, defined as the detection of some new information that is offered regarding previous knowledge (Ghosal et al.,, 2022), and cultural novelty as the subjective perception of cultural differences (Wilczewski et al.,, 2023). To establish this relationship, we first proposed a new dataset GloablFusion composed of cooking recipes linked to a specific country and variations adapted by different communities. Building a relevant knowledge space based on products whose textual description represents proxies for cultural aspects allowed us to build metrics that make it possible to interpret textual divergences as an indicator of cultural novelty and, thus, cultural differences. This framework exhibits the possibility of paralleling these two notions and shows the potential of natural language processing methods to represent this complex notion. In particular, we show that novelty metrics based on distribution divergence are efficient tools to measure cultural distance between the description of cultural products. Despite having relatively low R-square scores, we still observe significant and positive correlations that prove the utility of novelty detection techniques in language at least captures a small portion of such cultural differences. The gap is also easily understandable since these metrics rely on the partial descriptions of cultural products that do not embed the full spectrum of what constitutes cultural distances. However, this study demonstrates that a link exists and that some dimensions of language can be used to give a specific view of cultural differences.
@@
Another interesting point that increases the relevance of our metrics to capture novelty in such cultural products is their relation with crucial recipe aspects that also embed the notion of cultural novelty. In fact, innovation in food processing is perceived as a modification of ingredients, food preparation, its shape, its taste, or its packaging (Guerrero et al.,, 2009). The link emphasized between Surprise and, in a smaller portion, Difference metrics to the ratio of new ingredients and the Uniqueness to recipe lexical diversity describe some of these properties. It also highlights part of the limitation of our model, which is to capture the entirety of perceived innovation, since the textual information in our dataset cannot capture notions of shape, size, and taste.
@@
We also discuss the case of text representation and divergence metrics as tools for measuring novelty. We deliberately chose not to enhance text representation through methods like topic modeling or contextual embeddings derived from recent LLMs. While we acknowledge that such techniques could improve novelty modeling, they would complicate the interpretation of our results. We also recognize the limitation of divergences-based approaches for novelty detection since such divergences could stem from complex phenomena beyond novelty. However, our choices become evident in an evaluation context, such as assessing the quality of LLM-generated content. Suppose distribution divergence techniques are to be used as metrics to evaluate LLMs’ capacity for generating culturally aware novelty. In that case, it is crucial to ensure that the target of evaluation is not embedded within the measurement system itself (Strathern,, 1997). Using advanced text encoding approaches could introduce confounding factors, making it difficult to determine whether correlations with our cultural distance metrics arise from the language data or the encoding model. Finally, as LLMs training and fine-tuning rely on entropy-based learning, similar to the divergence measures employed in this study, a fine-grained understanding of how language distributions evolve across cultural contexts could offer valuable insights. Such research could inform strategies to fine-tune LLMs, reducing their inherent bias toward the dominant Anglophone culture (Tao et al.,, 2024).
@@
As the field of AI continues to expand, incorporating theoretical frameworks from social sciences becomes increasingly essential to address the ethical and cultural implications of AI tools. This aligns with recent calls in the literature to integrate sociological and management theories into AI research, especially for LLMs, to design systems and metrics better aligned with human values and societal needs (Liu et al.,, 2023). This study offers a concrete contribution in that direction by bridging the gap between artificial intelligence research and theoretical frameworks from sociology and management. By applying several transformations of the Jensen-Shannon divergence to measure the different kinds of novelty of the adaptation of cultural products in various countries and communities, we demonstrate how computational tools can integrate sociocultural insights to address complex questions about language, culture, and creativity. This interdisciplinary approach is critical for developing metrics that are not only robust but also contextually relevant in evaluating artificial intelligence systems, particularly as they aim to produce content that reflects diverse cultural perspectives.
@@ 7 Limitations and Future research avenues
A key limitation of our work is its reliance on a single dataset centered on one type of cultural product, justified by the absence of datasets dedicated to cultural novelty. This limitation is exacerbated since our novelty metrics only reflect a cultural dimension through the precise and strong cultural dimension of the knowledge space. However, developing GlobalFusion involved substantial manual effort to extract country-related recipes and balance country representation, diversity, and data sufficiency. Moreover, our approach introduced potential biases such as irrelevant recipe variations, popularity bias limiting underrepresented regions such as Africa or Oceania, cultural misappropriations, and the use of exclusively anglophone sources. We provide more details of these limitations in appendix A.1. To mitigate these issues, we intend to pursue this line of work by extending our work to other cultural products such as movie synopsis or literature book published and to take interest in database in different languages.
This research opens avenues for exploring alternative divergence representations, such as multi-prototypical and conceptual language representations, or entropy ranking-based approaches, to capture new dimensions of cultural novelty. Future work will also examine how different LLMs encode cultural variations and cultural biases in their embeddings, and assess their ability to generate coherent variations of cultural products, providing valuable insights into their representation of cultural novelty and diversity.
@@ Appendix A Dataset description
In this section, we introduce further information on our dataset. We created GlobalFusion our dataset by extracting around 130K recipes from recipeNLG (Bień et al.,, 2020), where we could identify the country or nationality in the title associated with the recipe. To extract the 500 dishes in our dataset, we followed a selection protocol based on the following conditions to identify the proper recipes for our analyses.
@@
Figure 6(a) introduces the total country repartition in the dataset. We then selected 500 recipes associated with a country to constitute our knowledge bases and to identify variations. Figure 6(b) presents the country repartition of these 500 recipes. The most represented countries in our dataset are Mexico, Italy, Greece, the United States, and Germany. The most represented countries in our 500 recipes are Mexico, Italy, Greece, the United States, and France.
Figure 6: Country repartition in GlobalFusion
@@
A.1 Dataset Limitations
Despite controlling manually and curating many recipe names, we still have undesirable recipe variations. For example, when collecting pancake recipes with also collected recipes such as pancakes with fried chicken, they do not correspond to the original cultural product and thus add novelty not related to cultural variations. The second limitation stems from a popularity bias; since we need enough recipes from a country to constitute a relevant knowledge base, we struggled to include a full diversity of countries from Africa, South America, or Oceania, limiting the broad scope of our analysis. A third limitations comes from certain countries’ tendency to appropriate cultural markers, such as countries’ names, in their culinary traditions and promote their cuisine. We can cite examples such as the "Indian Curry," a British dish, or the "French Silk Pie", a US one. Although we manually corrected these identified misappropriations, it indicates the possible presence of noise in our dataset. Finally, our last limitation comes from the language used. In fact, all our recipes are extracted from anglophone websites, creating a bias in the cultural representation of recipes and potentially affecting our analyses.
@@ B.2 Complementary results
$\begin{split}CulturalDistance=\beta_{0}+\beta_{1}*Newness+\beta_{2}*Uniqueness+\beta_{3}*Difference+\\ \beta_{4}*Surprise+\beta_{5}*NewIngredients+\beta_{6}*LexicalDiversity+\beta_{7}*LengthRatio\end{split}$ (9)
Table 5 presents our model’s detailed results.
@@
We also ran some complementary analyses to estimate the marginal contribution of each variable to the different cultural distances analyzed, compare them with the full model that includes the control variables, and better explain the observed Pearson correlations. The table 6 presents these results. We observe differences between these marginal contributions and the effect in the full model; knowing the correlation between our metrics and some metadata, such as the number of new ingredients, we also performed mediation analyses of this marginal contribution to have information on their direct effect on cultural distances. The table presents 7 these results. In this table, we do not present the mediation effects when the variable is already not significant in any way to the effect on cultural distances.

### v2

@@ 1 Introduction
In the context of NLP and AI, the emergence of generative large language models (LLMs) has spurred research into cultural diversity and awareness, primarily through external tasks involving prompting (Bhatt and Diaz,, 2024; Li et al.,, 2024). Among such approaches, researchers try to measure LLMs personality traits, values and cultural perspective through self-assessment tests and controllable prompts (Kovač et al.,, 2023). Recent work using the WVS Inglehart–Welzel cultural map (,, WVS) proposed interpretable cultural distance metrics to evaluate LLM outputs (Tao et al.,, 2024). While this approach highlights cultural representation differences, its reliance on specific questionnaires limits generalizability across tasks and datasets. In this paper, we propose five new cultural novelty detection metrics based on probability distribution divergences, an area of interest in out-of-distribution detection (Yang et al.,, 2024), which exhibits promising properties, such as entropic uncertainty, when applied to LLMs (Ahmadian et al.,, 2024). In sociology and management, divergence measures have been widely used to analyze cultural evolution, legal systems, societal norms (Klingenstein et al.,, 2014), and community values (Gallagher et al.,, 2018). This work bridges the gap between cultural novelty and distance in sociology and the challenges of cultural representation and diversity in AI. Leveraging divergence-based novelty detection, we propose new metrics to measure cultural differences through text analysis. By adapting these well-established textual novelty measures to the cultural domain, we aim to demonstrate their potential to meaningfully capture and generalize cultural differences across contexts, focusing on cultural proxies like food. Our contributions are as follows:
@@ 4.1 Dataset
To build our dataset on cultural novelty, we used RecipeNLG (Bień et al.,, 2020). The dataset notably comprises the title, a recipe description, and a list of ingredients of more than 2M recipes. To extract our culturally aware knowledge bases and variations, we kept the recipes where we could identify a country in the title. We ended up with 135 385 recipes from 173 different countries. Then, we defined product names as typical dishes, for example, "Lasagna," "Couscous," or "Beef Curry," which were associated with a country as our cultural variable. We have manually selected the dishes based on several conditions: (1) the dish must include enough recipe variation for evaluation; (2) the dishes should present geographical and ingredient diversity; (3) the dish must have enough ingredients, should be a dish on its own, and should not follow normative constraints. Using the recipe title to match the proper dishes, we extracted each cultural product’s associated recipe description to establish our knowledge space. We also identified variations of these dishes as recipe descriptions from different countries. Figure 2 presents an illustration of an image of recipes extracted with this approach for the couscous recipes, where the knowledge space is composed of 167 Moroccan recipes and 426 cultural variations coming from 35 different countries of all regions of the world.
@@
We also disclose two couscous recipes, one from Morocco and one from Thailand, to provide an example of recipe preparation and the type of semantic variation we will study with our metrics.
Moroccan Couscous Jamaican Chicken Couscous
Sprinkle the lamb with Moroccan spice. Heat the oil in a pan, add the lamb fillets and cook until the lamb is cooked as desired. Remove the lamb to a warm serving plate. Add the onion, garlic, red pepper and carrot to the pan and cook for 2-3 minutes. Add the tomatoes and juice, and the Baby Peas, stirring to combine. Cover and cook over a low heat for 10 minutes. Spoon the sauce onto the prepared couscous. Cut the lamb into slices and serve on top of the sauce. In a large saucepan, combine pineapple and 1/2 teaspoon salt; bring to a boil. Stir in couscous. Remove from heat; let stand, covered, 5 minutes or until liquid is absorbed. Fluff with a fork. Meanwhile, in a shallow bowl, mix flour and 2 tablespoons cilantro. Toss chicken with jerk seasoning and remaining salt. Add to flour mixture, a few pieces at a time, and toss to coat lightly; shake off any excess. In a large skillet, heat 1 tablespoon oil. Add a third of the chicken; cook 1-2 minutes on each side or until no longer pink. Repeat twice with remaining oil and chicken. Serve with couscous. If desired, sprinkle with cilantro.
Table 1: Two culturally distinct couscous recipes from our dataset, illustrating the adaptation of couscous across cultural contexts.
We identified 500 dishes, associated with 54 various countries from different geographical regions, forming our pairs of cultural products (product, country). Figure 3 presents the detailed repartition of countries of our selected dishes.
Figure 3: Country repartition of recipes in GlobalFusion.
We used the title to match a recipe and extracted the description to create our knowledge base. Each dish’s knowledge space comprises an average of 60.98 recipe descriptions and is paired with 208.96 recipe descriptions of new variants coming from 19.28 different countries, for a total of 104,710 recipes variations. The specific protocol for data collection, ensuring respect for several conditions presented and detailed statistics on the dataset are provided in the appendix A.
@@ 4.2 Experiment and Evaluation
  • We use the Inglehart–Welzel Cultural Map , (WVS) provided by the World Value Survey. This global cultural map scores society on two dimensions: (1) Traditional values versus Secular-rational values and (2) Survival values versus Self-expression values. Following the work in Tao et al., (2024), this two-dimensional measurement of culture allows us to estimate the Euclidean distance between countries studied.
  • We use the linguistic distance dataset [^3] http://dow.net.au/?page_id=32. We used the data collected from 2015 which followed the research protocol from Gordon Jr, (2005). This distance represents the distance between the two closest major languages for each pair of countries.
@@
The dataset and the code are available on our Github page [^5] https://github.com/fcarichon/Cultural_Novelty.
@@ 5 Results
Figure 5: Correlations between our 5 novelty metrics and 4 distinct cultural distances. The bars’ height highlights the strength of the coefficient, while the p-value is indicated above each bar.
@@ 6 Discussion
The present study took an interest in the link between textual novelty detection, defined as the detection of some new information that is offered regarding previous knowledge Ghosal et al., (2022), and cultural novelty as the subjective perception of cultural differences Wilczewski et al., (2023). To establish this relationship, we first proposed a new dataset GlobalFusion composed of cooking recipes linked to a specific country and variations adapted by different communities. Building a relevant knowledge space based on products whose textual description represents proxies for cultural aspects allowed us to build metrics that make it possible to interpret textual divergences as an indicator of cultural novelty and, thus, cultural differences. This framework exhibits the possibility of paralleling these two notions and shows the potential of natural language processing methods to represent this complex notion. In particular, we show that well-established general novelty metrics based on distribution divergence are efficient tools to measure cultural distance between the description of cultural products. Despite having relatively low R-square scores, we still observe significant and positive correlations that prove the utility of novelty detection techniques in language at least captures a small portion of such cultural differences. The gap is also easily understandable since these metrics rely on the partial descriptions of cultural products that do not embed the full spectrum of what constitutes cultural distances. However, this study demonstrates that a link exists and that some dimensions of language can be used to give a specific view of cultural differences.
@@
Another interesting point that increases the relevance of our metrics to capture novelty in such cultural products is their relation with crucial recipe aspects that also embed the notion of cultural novelty. In fact, innovation in food processing is perceived as a modification of ingredients, food preparation, its shape, its taste, or its packaging Guerrero et al., (2009). The link emphasized between Surprise and, in a smaller portion, Difference metrics to the ratio of new ingredients and the Uniqueness to recipe lexical diversity describe some of these properties. It also highlights part of the limitation of our model, which is to capture the entirety of perceived innovation, since the textual information in our dataset cannot capture notions of shape, size, and taste.
@@
We also discuss the case of text representation and divergence metrics as tools for measuring novelty. We deliberately chose not to enhance text representation through methods like topic modeling or contextual embeddings derived from recent LLMs. While we acknowledge that such techniques could improve novelty modeling, they would complicate the interpretation of our results. We also recognize the limitation of divergences-based approaches for novelty detection since such divergences could stem from complex phenomena beyond novelty. However, our choices become evident in an evaluation context, such as assessing the quality of LLM-generated content. Suppose distribution divergence techniques are to be used as metrics to evaluate LLMs’ capacity for generating culturally aware novelty. In that case, it is crucial to ensure that the target of evaluation is not embedded within the measurement system itself Strathern, (1997). Using advanced text encoding approaches could introduce confounding factors, making it difficult to determine whether correlations with our cultural distance metrics arise from the language data or the encoding model. Finally, as LLMs training and fine-tuning rely on entropy-based learning, similar to the divergence measures employed in this study, a fine-grained understanding of how language distributions evolve across cultural contexts could offer valuable insights. Such research could inform strategies to fine-tune LLMs, reducing their inherent bias toward the dominant Anglophone culture Tao et al., (2024).
@@
As the field of AI continues to expand, incorporating theoretical frameworks from social sciences becomes increasingly essential to address the ethical and cultural implications of AI tools. This aligns with recent calls in the literature to integrate sociological and management theories into AI research, especially for LLMs, to design systems and metrics better aligned with human values and societal needs Liu et al., (2023). This study offers a concrete contribution in that direction by bridging the gap between artificial intelligence research and theoretical frameworks from sociology and management. By applying several transformations of the Jensen-Shannon divergence to measure the different kinds of novelty of the adaptation of cultural products in various countries and communities, we demonstrate how computational tools can integrate sociocultural insights to address complex questions about language, culture, and creativity. This interdisciplinary approach is critical for developing metrics that are not only robust but also contextually relevant in evaluating artificial intelligence systems, particularly as they aim to produce content that reflects diverse cultural perspectives.
@@ 7 Limitations and Future research avenues
A key limitation of our work is its reliance on a single dataset focused on one type of cultural product due to the lack of datasets dedicated to cultural novelty. This is compounded by our novelty metrics reflecting culture through the precise cultural representation in the knowledge space. However, developing GlobalFusion involved a substantial manual effort to extract country-related recipes, balance representation and diversity, and ensure data sufficiency. It also required minimizing noise and cultural misappropriation in title matching and addressing biases from varying writing norms. While literature, news, and movie synopses are promising directions for future work, assigning a cultural identity to them is more complex than for recipes, where users explicitly self-identify the country of origin. For instance, defining a movie’s cultural origin, whether by its producer, director, actors, or dominant themes, poses methodological challenges requiring careful consideration to avoid biases.
A second limitation stems from our selection strategy, which introduces several cultural representation biases. First, we struggled to include a complete diversity of countries from Africa, South America, or Oceania due to RecipeNLG’s limited coverage, narrowing the scope of our analysis. Additionally, certain countries use appropriate cultural markers, such as countries’ names, in their culinary traditions; for example, "Indian curry" is a British dish. Although we manually corrected identified cases, other misattributions may remain. Finally, our last limitation comes from the language used. All our recipes are sourced from anglophone websites, creating linguistic and cultural representation biases. Using anglophone data for foreign recipes introduces alteration preparation descriptions. However, multilingual recipe datasets are scarce, and translating recipes from original languages could introduce biases of modifying the phraseology of recipes or lacking grounding from these specific cultures Hershcovich et al., (2022); Cao et al., (2024).
This research opens avenues for exploring alternative divergence representations, such as multi-prototypical and conceptual language representations or entropy ranking-based approaches, to capture new dimensions of cultural novelty. Moreover, our metrics could be applied to estimate the propagation of novel recipes across geographical boundaries using temporal metadata. They also open avenues in the classic NLP field and LLM analyses. First, traditional measures like TF-IDF and cosine similarity may enhance these metrics; it could also help understand cultural representation in word embeddings. For LLMs, prompting models to generate either traditional Moroccan or Venezuelan couscous would allow for measuring how well they reflect culturally grounded adaptations. Finally, since LLM fine-tuning often leverages relative entropy Ahmadian et al., (2024), our metrics align naturally with such adaptation tasks and could be integrated into model evaluation or mitigation processes.
@@ Appendix A Dataset description
In this section, we introduce further information on our dataset. We created GlobalFusion our dataset by extracting around 130K recipes from recipeNLG Bień et al., (2020), where we could identify the country or nationality in the title associated with the recipe. To extract the 500 dishes in our dataset, we followed a selection protocol based on the following conditions to identify the proper recipes for our analyses.
@@
Figure 7 introduces the total country repartition in the dataset. The countries most represented in our dataset are Mexico, Italy, Greece, the United States, and Germany. We then selected 500 recipes associated with a country to constitute our knowledge bases and to identify variations. The countries most represented in our 500 recipes are Mexico, Italy, Greece, the United States, and France.
Figure 7: Total Country repartition in GlobalFusion.
@@ B.2 Complementary results
$\begin{split}CulturalDistance=\beta_{0}+\beta_{1}*Newness+\beta_{2}*Uniqueness+\\ \beta_{3}*Difference+\beta_{4}*Surprise+\beta_{5}*NewIngredients\\ +\beta_{6}*LexicalDiversity+\beta_{7}*LengthRatio\end{split}$ (9)
Table 6 presents the detailed results of our model.
@@
We also ran some complementary analyses to estimate the marginal contribution of each variable to the different cultural distances analyzed, compare them with the full model that includes the control variables, and better explain the observed Pearson correlations. The table 7 presents these results. We observe differences between these marginal contributions and the effect in the full model; knowing the correlation between our metrics and some metadata, such as the number of new ingredients, we also performed mediation analyses of this marginal contribution to have information on their direct effect on cultural distances. The table presents 8 these results. In this table, we do not present the mediation effects when the variable is already not significant in any way to the effect on cultural distances.